OpenStudy (anonymous):

A system of equations is shown below: y = 5x + 3 y = 4x + 6 What is the solution to the system of equations? (−3, 18) (−3, −18) (3, −18) (3, 18)

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

@EclipsedStar

OpenStudy (anonymous):

There are multiple ways to solve your questions, one way is to plug, and play your answer choices. \((x, y) \rightarrow (3, 18)\) Try, and see if that works for both of your equations.

OpenStudy (anonymous):

yes correct yay

OpenStudy (anonymous):

Line A is represented by the equation given below: x + y = 2 What is most likely the equation for line B, so that the set of equations has infinitely many solutions? 2x + 2y = 2 2x + 2y = 4 2x + y = 2 x + y = 4

OpenStudy (anonymous):

what is it @Algorithmic